OpenStudy (anonymous):

mitosuki Neophyte solve √x - √5 + 7 = x (The radical is suppose to go over the whole x - 5) a.x = –6 b.x = –9 c.x = –6 and x = –9 d.None of the above

OpenStudy (jamesj):

Write \[\sqrt{x-5} = x - 7\] and now square both sides, giving you ....

myininaya (myininaya):

\[\sqrt{x}-\sqrt{5} \neq \sqrt{x-5}\]

myininaya (myininaya):

but i assume you meant \[\sqrt{x-5}\]

OpenStudy (anonymous):

yes that is what I meant

myininaya (myininaya):

so what happens after completing the step james mentioned

OpenStudy (anonymous):

Ok the answer I'm getting is x = 6 and x = 9 but only x=9 works. But none of the answers are there. So it would just be none of the above right? I want to make I'm not doing anything wrong.

myininaya (myininaya):

\[(\sqrt{x-5})^2=(x-7)^2\] is this what he means by sqaring both sides? if so, then if you square the square root of something what do you get? that something right? so we have \[x-5=(x-7)^2\]

myininaya (myininaya):

\[x-5=(x-7)(x-7)\] \[x-5=x(x-7)-7(x-7)\] \[x-5=x^2-7x-7x+49\] \[x-5=x^2-14x+49\] put everything on one side by subtracting x on both sides and by adding 5 on both sides \[0=x^2-14x-x+49+5\]

myininaya (myininaya):

\[0=x^2-15x+54\] so we need to find two factors of 54 that have product 54 and have sum -15 54=-6(-9) -15=-6-9 0=(x-6)(x-9) x=6 or x=9 or is it neither? check both answers to see the oringral equation was \[\sqrt{x-5}+7=x\] what happens when you plug in 6 1+7=6 noy true 6 is not a solution what happens when you plug in 9 2+7=9 is a true statement x=9 is only solution

OpenStudy (jamesj):

sigh. I hate it when I can't add.

myininaya (myininaya):

the only mistake i think i see is i don't know how to spell oringal

myininaya (myininaya):

adding sucks

OpenStudy (anonymous):

lol well thank you guys very much for the help :)

myininaya (myininaya):

oops another mistake not noy but not lol

myininaya (myininaya):

1+7=6 not true*

OpenStudy (anonymous):

Oki doki thank you

myininaya (myininaya):

np

myininaya (myininaya):

oh yeah good job mitosuki this is what you said you got right? you deserve a medal :)

OpenStudy (anonymous):

haha yes I did thank you again C :

OpenStudy (anonymous):

*C: